# Adaptive Mix for Semi-Supervised Medical Image Segmentation

### v1

@@ 4.6 Analysis of Adaptive Mix Samples
AdaMix provides perturbed image examples with gradually increasing sample complexity, facilitating the effective consistency regularization on unlabeled data. As shown in Fig.5, taking four cases from the ACDC dataset as an example, we illustrate the AdaMix procedure during training. Overall, we can observe that high-confidence regions typically appear in the background, while low-confidence ones tend to appear at the boundaries of the segmentation targets. In medical image segmentation tasks, the background region contains simple patterns, while the object boundaries are considered complex ones (Yu et al., 2019; Shen et al., 2023). Deep neural networks learn simple patterns first (Arpit et al., 2017). In the initial stage of training, considering the model’s inferior capability, AdaMix provides perturbed samples with more high-confidence regions (i.e., more background regions), from which the model can generate relatively accurate predictions. As training progresses, the model’s performance improves; AdaMix gradually incorporates low-confidence patches that generally appear at challenging target boundaries into the perturbed images, providing more complex samples for training; forcing the model to focus on these regions enhances its discriminative capability in the later training stages. In light of the above, AdaMix can adaptively adjust the perturbation strength based on the model’s learning state, thereby achieving more efficient consistency learning.
4.7 Qualitative Results
As shown in Fig. 6, we visualize some segmentation examples on the ACDC, ISIC, and LA datasets. Specifically, the supervised baseline is prone to under-segment some lesion regions due to the limited number of labeled data. FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021) (all using CutMix (Yun et al., 2019) as the strong perturbation) incorrectly classify many boundary pixels, mainly due to noise introduced by random perturbations during training. Similarly, UCMT (Shen et al., 2023) and ABD (Chi et al., 2024), which use UMix (Shen et al., 2023) as the strong perturbation, also exhibit inferior performance on segmentation boundaries. These results can be attributed to the limited effectiveness of consistency regularization. In contrast, our AdaMix-ST exhibits smoother and clearer boundaries compared to other methods. The finer segmentation results can be ascribed to the AdaMix algorithm which adaptively generates perturbed images in a self-paced learning manner.
@@ 5 Conclusion
This paper explores adaptive mix-up techniques for semi-supervised medical image segmentation. We propose a novel Adaptive Mix algorithm (AdaMix) for strong-weak pseudo supervision in semi-supervised learning. AdaMix features a self-paced curriculum that initially provides relatively simple mixed samples and then gradually increases the difficulty of the mixed images by adaptively controlling the mix-up strength based on the model’s learning state. AdaMix can be seamlessly integrated as a plug-and-play module into self-training, mean-teacher, and co-training semi-supervised learning paradigms, achieving superior performance. Experiments on three public medical image datasets validate the effectiveness of the proposed method and demonstrate its superiority over state-of-the-art approaches. Moreover, we found that replacing high-confidence regions with low-confidence ones, though seemingly counterintuitive, is effective for consistency regularization. Since low-confidence regions often appear at segmentation boundaries, forcing the model to learn from these areas helps enhance its discriminative capability. We also found that, in strong-weak pseudo supervision for semi-supervised learning, perturbation strategies are more essential than learning paradigms such as self-training, mean-teacher, and co-training.
@@
Zhiqiang Shen: Writing – original draft, Methodology, Visualization, Validation, Data curation, Conceptualization. Peng Cao: Writing – review & editing, Supervision, Project administration, Investigation, Funding acquisition, Formal analysis, Conceptualization. Junming Su: Data curation, Visualization, Validation. Jinzhu Yang: Investigation, Funding acquisition. Osmar Zaiane: Writing – review & editing, Formal analysis.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62076059) and the Science and Technology Joint Project of Liaoning Province (Grant No. 2023JH2/101700367, ZX20240193).

### v2

@@ 4.4.6 Analysis of Adaptive Mix Samples
AdaMix provides perturbed image examples with gradually increasing sample complexity, facilitating the effective consistency regularization on unlabeled data. As shown in Fig. 9, taking four cases from the ACDC dataset as an example, we illustrate the AdaMix procedure during training. Overall, we can observe that high-confidence regions typically appear in the background, while low-confidence ones tend to appear at the boundaries of the segmentation targets. In medical image segmentation tasks, the background region contains simple patterns, while the object boundaries are considered complex ones (Yu et al., 2019; Shen et al., 2023). Deep neural networks learn simple patterns first (Arpit et al., 2017). In the initial stage of training, considering the model’s inferior capability, AdaMix provides perturbed samples with more high-confidence regions (i.e., more background regions), from which the model can generate relatively accurate predictions. As training progresses, the model’s performance improves; AdaMix gradually incorporates low-confidence patches that occur at challenging target boundaries into the perturbed images, providing more complex samples for training; forcing the model to focus on these regions enhances its discriminative capability in the later training stages. In light of the above, AdaMix can adaptively adjust the perturbation strength based on the model’s learning state, achieving more efficient consistency regularization.
@@ 5 Conclusion
This paper explores adaptive mix-up techniques for semi-supervised medical image segmentation. We propose a novel Adaptive Mix algorithm (AdaMix) that features a self-paced curriculum, which initially provides relatively simple mixed samples and then gradually increases the difficulty of the mixed images by adaptively controlling the mix-up strength based on the model’s learning state. AdaMix can be seamlessly integrated as a plug-and-play module into self-training, mean-teacher, and co-training semi-supervised learning paradigms, achieving superior performance. Experiments on three public medical image datasets validate the effectiveness of the proposed method and demonstrate its superiority over state-of-the-art approaches. Moreover, we reveal that incorporating more high-confidence patches into images generates easier samples, which benefits the segmentation model during the initial training stage; in contrast, replacing high-confidence regions with low-confidence ones produces harder examples that are effective for the later training stage. Since low-confidence regions often appear in segmentation boundaries, forcing the model to learn from these regions helps enhance its discriminative capability. Our analysis also demonstrates that, in strong-weak pseudo supervision for semi-supervised learning, perturbation strategies are more critical than learning paradigms.
@@
Zhiqiang Shen: Writing – original draft, Methodology, Visualization, Validation, Data curation, Conceptualization. Peng Cao: Writing – review & editing, Supervision, Project administration, Investigation, Funding acquisition, Formal analysis, Conceptualization. Junming Su: Data curation, Visualization, Validation. Jinzhu Yang: Investigation, Funding acquisition. Osmar R. Zaiane: Writing – review & editing, Formal analysis.
@@ Acknowledgments
This research was supported by the National Natural Science Foundation of China (No.62076059), the Science and Technology Joint Project of Liaoning Province (2023JH2/101700367), the Fundamental Research Funds for the Central Universities (N2424010-7), and the China Scholarship Council (202406080040).